# Mitigating Hallucinations in Vision-Language Models through Image-Guided Head Suppression

### v1

@@ 4.2 CHAIR Evaluation
Mitigating Repetition from Head Pruning: In our experiments, we observed that SPIN occasionally leads to repetitive outputs for Minigpt4. While SPIN effectively reduces hallucinations and maintains F1 scores close to the baseline, it increases the length of generated sequences by up to 3.8. To mitigate this, we leverage the repetition penalty during decoding, as shown in Table 3. While a high value can negatively impact F1, a repetition penalty of 1.1 alleviates repetitions without compromising overall output quality, as evidenced by caption length and F1 performance.
Method | Rep. | Scale | CHAIRS | CHAIRI | F1 | Caption |
Penalty | Factor | () | () | () | Length | |
Baseline | 1.0 | - | 31.4 | 11.1 | 70.6 | 84.1 |
SPIN (r=0.15, Layers 124) | 1.0 | 0.0 | 19.6 | 6.8 | 68.2 | 321.3 |
1.05 | 0.0 | 19.2 | 7.7 | 68.9 | 100.1 | |
1.08 | 0.0 | 20.2 | 7.8 | 68.1 | 57.4 | |
1.1 | 0.0 | 19.2 | 7.8 | 67.5 | 42.9 | |
SPIN (r=0.18, Layers 116) | 1.0 | 0.0 | 21.0 | 6.2 | 68.8 | 271.6 |
1.1 | 0.0 | 17.4 | 8.3 | 67.7 | 43.0 | |
1.1 | 0.05 | 17.6 | 8.4 | 68.4 | 44.7 | |
1.1 | 0.08 | 21.6 | 9.2 | 69.4 | 46.2 |
@@ 4.3 POPE Evaluation
Method: Polling-based Object Probing Evaluation (POPE) Li et al. (2023) is used for assessing hallucinations in VQA tasks by querying models with “Is there a <object> in the image?”. The objects are drawn from three splits: random (any dataset object), popular (most frequent objects), and adversarial (closely related but misleading objects). We evaluate on 500 randomly selected COCO 2014 Lin et al. (2014) validation images, with six questions per image from each split. We present results for multi-turn POPE evaluation Liu et al. (2024c), where earlier responses are appended to the input context, increasing the context length and amplifying the chances of image neglect.
Evaluation on greedy decoding: Table 2 reports POPE results for LLaVA (7B and 13B), Minigpt4 and Shikra across all three splits using greedy decoding. SPIN consistently improves accuracy and F1 across all three splits for LLaVA and Shikra models, achieving gains of up to 3.4% in accuracy and 3.7% in F1 over the baseline. DAMRO performs similarly or worse than baseline for greedy decoding. SPIN consistently outperforms PAI in both accuracy and F1 for LLaVA and Shikra, achieving up to a 2.6% gain in both metrics. For MiniGPT-4, SPIN attains the highest F1 score, while PAI achieves slightly better accuracy on the popular and adversarial splits.
Decoding | Mode | Accuracy | F1 |
Beam Search | Baseline | 83.58 | 81.14 |
OPERA | 83.67 | 81.29 | |
PAI | 84.52 | 82.80 | |
SPIN | 86.15 | 84.93 | |
Nucleus | Baseline | 82.97 | 81.70 |
VCD | 84.42 | 83.45 | |
PAI | 84.96 | 83.99 | |
DAMRO | 85.69 | 84.28 | |
SPIN | 82.49 | 81.83 |
Evaluation on other decoding strategies: In Table 4, we report overall accuracy and F1 scores for multi-turn POPE evaluation using beam-search and nucleus sampling decoding. We observe that SPIN yields the best results for beam-search, whereas DAMRO performs the best for nucleus sampling. We observe that SPIN performs more effectively with greedy decoding and beam-search, but is less effective under nucleus sampling. This may be attributed to the inherent randomness in sampling-based methods, which introduce greater variability in token selection. As a result, even after suppressing hallucination-prone heads, the stochastic nature of nucleus sampling can still lead to hallucinated outputs, limiting the impact of structural interventions like head suppression.
Method | LLaVA-1.5(7B) | LLaVA-1.5(13B) | Qwen-VL | |||
A | D | A | D | A | D | |
Baseline | 6.50 | 6.83 | 6.53 | 6.71 | 6.56 | 6.87 |
SPIN | 7.19 | 6.12 | 7.03 | 6.18 | 7.08 | 6.22 |
@@ 4.4 MMHal-Bench Evaluation
Evaluation Results: The final evaluation report including per-category scores for each model, with the overall performance measured as the average across all categories, is presented in Figure 4. We observe that SPIN improves overall scores over the baseline for all three models: LLaVa-1.5, Minigpt4 and Shikra, with the maximum improvements observed in holistic description and other categories.
Layers | CS | CI | F1 | |
132 | 0.05 | 17.4 | 5.8 | 74.6 |
132 | 0.10 | 22.8 | 7.4 | 74.1 |
132 | 0.15 | 20.8 | 6.4 | 73.8 |
116 | 0.05 | 18.2 | 6.2 | 73.7 |
120 | 0.05 | 18.4 | 6.3 | 74.1 |
124 | 0.05 | 19.2 | 6.5 | 73.8 |
432 | 0.05 | 42.0 | 11.4 | 78.2 |
832 | 0.05 | 42.4 | 11.1 | 78.8 |
Layers | CS | CI | F1 | |
132 | 0.05 | 45.8 | 12.0 | 78.2 |
132 | 0.10 | 29.6 | 8.0 | 78.9 |
132 | 0.15 | 41.6 | 10.3 | 80.2 |
132 | 0.20 | 36.0 | 10.0 | 79.1 |
116 | 0.10 | 30.6 | 8.3 | 79.6 |
120 | 0.10 | 29.2 | 7.9 | 79.1 |
124 | 0.10 | 28.0 | 9.2 | 79.3 |
Layers | CS | CI | F1 | |
132 | 0.05 | 44.4 | 11.7 | 77.6 |
132 | 0.10 | 39.8 | 10.9 | 77.1 |
132 | 0.15 | 41.6 | 11.7 | 77.5 |
132 | 0.20 | 39.6 | 11.2 | 76.9 |
132 | 0.25 | 36.4 | 9.7 | 77.3 |
132 | 0.30 | 26.0 | 9.1 | 75.6 |
120 | 0.30 | 27.6 | 7.9 | 76.1 |
124 | 0.30 | 28.4 | 8.3 | 76.2 |
832 | 0.30 | 32.4 | 9.4 | 76.3 |
@@ 4.5 GPT-4o Assisted Evaluation
We use GPT-4o to compare the responses of baseline and SPIN models using Greedy decoding on a scale of 1 to 10 using two metrics: (1) Accuracy (A): measuring if the description is accurate with respect to the image content, and (2) Detailedness (D): measuring the richness of necessary details in the responses. The prompt we used is given in the Appendix Table 10. It is designed to ensure that the sequential order in which the responses are presented does not affect the judgement. The evaluation model is prompted to identify objects misaligned with the image context, identifying any discrepancies in count, position, and colors of objects in the images. The results presented in Table 5 show that SPIN improves Accuracy up to 7%, while maintaining Detailedness close to baseline.
@@ 4.6 Throughput Estimation
To assess the efficiency of different hallucination mitigation strategies, we compare their throughput in Figure 6. Throughput is computed by measuring the total computational latency for generating tokens, and dividing the total number of generated tokens by this latency. SPIN achieves the highest throughput among SoTA approaches, while achieving throughput performance close to baseline.
@@ 4.7 Ablation Studies
We present the hyperparameter evaluations for SPIN in Table 6, where we first vary to obtain reduced hallucinations without a significant drop in F1. We obtain =0.05 for LLaVA (7B), 0.10 for LLaVA (13B), and 0.30 for Qwen-VL. Next, we evaluate different layer configurations to find hallucination-prone layers . While uniformly pruning across all layers works best for LLaVA (7B), initial layers are more hallucination-prone for LLaVA (13B) and Qwen-VL. Finally, for selected and layers for head suppression, we plot , and F1 for different in Figure 5. F1 scores show comparable values for different for Minigpt4 and LLaVa 13B models. However, for LLaVa 7B and Qwen-VL, complete pruning leads to a drop in F1, which can be effectively mitigated by using a higher .
@@ 5 Conclusions
We propose an inference-only solution for ease of integration and to accommodate scenarios with limited training resources. However, our approach could be enhanced by incorporating a trainable router for more adaptive head selection. Our method’s effectiveness is somewhat reduced when paired with sampling-based decoding strategies (e.g., nucleus sampling). We attribute this to the greater token selection variability inherent in these stochastic approaches (as further detailed in Section 4). Our method requires access to the model weights, which limits its applicability to open-source models. It cannot be applied to closed-source or API-based models, where direct access to weights is restricted.
@@
Appendix A Appendix
A.1 Baseline Hyperparameters
The hyperparameters and for SPIN are specified in each table. For existing methods, we use the best-performing configurations reported in their respective papers.
For PAI, we adopt the following hyperparameter settings: for LLaVA, we use , , and ; for MiniGPT4, , , and ; and for Shikra, , , and . Since the original PAI paper does not report results on Qwen-VL, we experimented with several configurations and found that reusing Shikra’s settings yielded the best performance.
For OPERA, we use the same configuration across all models: , , , and . For VCD, we use , , and apply noise steps for all models. For DAMRO, we select the top 10 tokens as outliers and set .
@@ A.2 Ablation Studies
Head Selection Strategy: We explore a variety of training-free head selection strategies, as summarized in Table 7. Query-based head selection proposed in Jin et al. (2024) is computed based on the L2 norm of the query tokens, whereas the key based selection is performed based on the L2 norm of the key tokens. In addition, we experiment with head selection guided by the overall attention (to both text and image tokens) as well as attention to the image tokens only. Among these strategies, selecting heads based solely on image attention proves to be the most effective for hallucination reduction, yielding the best CHAIR scores. Here, we use , completely pruning attention heads across all layers.
Selection | Supp. | CHAIRS | CHAIRI | F1 |
Strategy | Heads | () | () | () |
Query | 0.10 | 40.4 | 11.6 | 77.4 |
Query | 0.25 | 34.6 | 10.4 | 75.8 |
Key | 0.25 | 44.6 | 12.4 | 77.1 |
Attention | 0.15 | 20.4 | 8.0 | 73.4 |
Attention | 0.05 | 17.6 | 6.1 | 74.5 |
Image Attention | 0.05 | 17.4 | 5.8 | 74.6 |
@@ A.3 Head Mask Visualization
The head masks for POPE and CHAIR using 95% active heads are visualized in Figure 7. The pruned heads are more consistent for POPE, likely due to the fact that “Yes” or “No” answers are generated as the first token. In contrast, CHAIR generates longer sequences, causing tokens at each position to likely prune different heads.
Model | Layers | Accuracy | F1 | ||
LLaVA-1.5 (7B) | - | - | - | 85.69 | 84.28 |
116 | 0.15 | 0.0 | 87.84 | 87.13 | |
116 | 0.15 | 0.001 | 87.79 | 87.07 | |
132 | 0.20 | 0.08 | 88.19 | 87.42 | |
LLaVA-1.5 (13B) | - | - | - | 85.12 | 84.48 |
132 | 0.35 | 0.0 | 88.47 | 88.00 | |
124 | 0.35 | 0.0 | 88.40 | 88.01 | |
120 | 0.35 | 0.001 | 88.50 | 88.17 | |
Minigpt4 | - | - | - | 78.08 | 76.20 |
132 | 0.05 | 0.0 | 79.43 | 78.78 | |
124 | 0.05 | 0.08 | 79.37 | 78.61 | |
124 | 0.05 | 0.001 | 79.14 | 78.46 | |
Shikra | - | - | - | 78.27 | 78.89 |
116 | 0.20 | 0.0 | 79.20 | 79.40 | |
116 | 0.20 | 0.01 | 79.17 | 79.37 | |
116 | 0.20 | 0.05 | 79.19 | 79.45 |

### v2

@@ 4.2 CHAIR Evaluation
In our experiments, we observed that SPIN occasionally leads to repetitive outputs for MiniGPT-4. While SPIN effectively reduces hallucinations and maintains F1 scores close to the baseline, it increases the length of generated sequences by up to 3.8. To mitigate this, we leverage the repetition penalty during decoding, as shown in Table 3. While a high value can negatively impact F1, a repetition penalty of 1.1 alleviates repetitions without compromising overall output quality, as evidenced by caption length and F1 performance.
Method
Rep.
Scale
CHAIRS
CHAIRI
F1
Caption
Penalty
Factor
Length
Baseline
1.0
31.4
11.1
70.6
84.1
SPIN
(r=0.15,
Layers
124)
1.0
0.0
19.6
6.8
68.2
321.3
1.05
0.0
19.2
7.7
68.9
100.1
1.08
0.0
20.2
7.8
68.1
57.4
1.1
0.0
19.2
7.8
67.5
42.9
SPIN
(r=0.18,
Layers
116)
1.0
0.0
21.0
6.2
68.8
271.6
1.1
0.0
17.4
8.3
67.7
43.0
1.1
0.05
17.6
8.4
68.4
44.7
1.1
0.08
21.6
9.2
69.4
46.2
Table 3: Analysis using repetition penalty on MiniGPT-4
@@ 4.3 POPE Evaluation
Method: Polling-based Object Probing Evaluation (POPE) Li et al. (2023) is used for assessing hallucinations in VQA tasks by querying models with “Is there a <object> in the image?”. The objects are drawn from three splits: random (any dataset object), popular (most frequent objects), and adversarial (closely related but misleading objects). We evaluate on 500 randomly selected COCO 2014 Lin et al. (2014) validation images, with six questions per image from each split. We present results for multi-turn POPE evaluation Liu et al. (2024d), where earlier responses are appended to the input context, increasing the context length and amplifying the chances of image neglect.
Evaluation on greedy decoding:
Table 2 reports POPE results for LLaVA (7B and 13B), MiniGPT-4 and Shikra across all three splits using greedy decoding. SPIN consistently improves accuracy and F1 across all three splits for LLaVA and Shikra models, achieving gains of up to 3.4% in accuracy and 3.7% in F1 over the baseline. DAMRO performs similarly or worse than baseline for greedy decoding. SPIN consistently outperforms PAI in both accuracy and F1 for LLaVA and Shikra, achieving up to a 2.6% gain in both metrics. For MiniGPT-4, SPIN attains the highest F1 score, while PAI achieves slightly better accuracy on the popular and adversarial splits. We provide POPE evaluation results for LLaVA-Next and Qwen2.5-VL in the Appendix Table 13.
Decoding
Mode
Accuracy
F1
Beam
Search
Baseline
83.58
81.14
OPERA
83.67
81.29
PAI
84.52
82.80
SPIN
86.15
84.93
Nucleus
Baseline
82.97
81.70
VCD
84.42
83.45
PAI
84.96
83.99
DAMRO
85.69
84.28
SPIN
82.49
81.83
Table 4: Multi-turn POPE evaluation using Beam Search and Nucleus Sampling decoding for LLaVA-1.5 (7B).
Evaluation on other decoding strategies:
In Table 4, we report overall accuracy and F1 scores for multi-turn POPE evaluation using beam-search and nucleus sampling decoding. We observe that SPIN yields the best results for beam-search, whereas DAMRO performs the best for nucleus sampling.
We observe that SPIN performs more effectively with greedy decoding and beam-search, but is less effective under nucleus sampling. This may be attributed to the inherent randomness in sampling-based methods, which introduce greater variability in token selection. As a result, even after suppressing hallucination-prone heads, the stochastic nature of nucleus sampling can still lead to hallucinated outputs, limiting the impact of structural interventions like head suppression.
@@ 4.4 MMHal-Bench Evaluation
Evaluation Results: The final evaluation report including per-category scores for each model, with the overall performance measured as the average across all categories, is presented in Figure 4. We observe that SPIN improves overall scores over the baseline for all three models: LLaVA-1.5, MiniGPT-4 and Shikra, with the maximum improvements observed in holistic description and other categories. Note, we use the same values for Layers, , and as reported for CHAIR evaluation.
Model
Method
Perception
Cognition
Exis-
tence
Count
Posi-
tion
Color
Posters
Cele-
brity
Scene
Land-
mark
Art-
work
OCR
Total
Comm-
onsense
Num-
erical
Text
Code
Total
LLaVA-
1.5 (7B)
Baseline
190.0
108.3
96.7
135.0
140.8
134.7
153.5
140.0
75.3
92.5
1266.8
93.6
77.5
45.0
45.0
261.1
SPIN
190.0
113.3
101.7
135.0
142.9
137.4
151.8
138.5
76.0
107.5
1294.0
93.6
75.0
60.0
45.0
273.6
LLaVA-
1.5 (13B)
Baseline
185.0
110.0
100.0
135.0
146.6
121.8
161.3
160.3
87.0
117.5
1324.4
110.7
52.5
90.0
47.5
300.7
SPIN
185.0
110.0
100.0
135.0
144.2
121.5
162.0
158.8
90.8
117.5
1324.7
112.1
62.5
87.5
47.5
309.6
Qwen2.5-
VL (7B)
Baseline
180.0
120.0
148.3
190.0
169.4
117.9
139.5
123.3
141.0
192.5
1521.9
110.7
155.0
185.0
140.0
590.7
SPIN
185.0
125.0
148.3
190.0
173.1
135.6
137.5
131.3
134.0
177.5
1537.3
110.0
125.0
185.0
132.5
552.5
Table 5: Results on MME evaluation. SPIN improves performance across both perception and cognition subsets on all three models. A small degradation in cognitive performance is observed for Qwen2.5-VL.
Model
Mode
Art &Design
Business
Science
Health &Medicine
Human. &Social Sci.
Tech &Eng.
Overall
LLaVA-1.5 (7B)
Baseline
51.7
23.3
24.7
34.0
49.2
31.4
34.4
SPIN
50.0
24.0
25.3
35.3
47.5
31.4
34.4
LLaVA-1.5 (13B)
Baseline
51.7
23.3
28.0
39.3
52.5
32.4
36.6
SPIN
51.7
23.3
29.3
38.0
54.2
32.9
36.9
LLaVA-Next (7B)
Baseline
52.5
30.0
20.7
36.7
57.5
26.2
35.3
SPIN
52.5
29.3
20.7
34.0
57.5
29.5
35.6
Qwen2.5-VL (7B)
Baseline
68.3
40.0
40.0
55.3
67.5
39.5
49.9
SPIN
67.5
42.7
37.3
54.7
68.3
40.0
49.9
Table 6: Performance on MMMU validation set for evaluating general capability. SPIN maintains or slightly improves general multimodal reasoning capabilities across all evaluated LVLMs. Following the original MMMU Yue et al. (2024) evaluation protocol, the overall score is computed as a weighted average across all subtasks.
@@ 4.5 GPT-4o Assisted Evaluation
Method
LLaVA-1.5 (7B)
LLaVA-1.5 (13B)
Qwen-VL
A
D
A
D
A
D
Baseline
6.50
6.83
6.53
6.71
6.56
6.87
SPIN
7.19
6.12
7.03
6.18
7.08
6.22
Table 7: GPT-4o assisted evaluation showcasing Accuracy (A) and Detailedness (D) of the generated captions.
We use GPT-4o to compare the responses of baseline and SPIN models using Greedy decoding on a scale of 1 to 10 using two metrics: (1) Accuracy (A): measuring if the description is accurate with respect to the image content, and (2) Detailedness (D): measuring the richness of necessary details in the responses. The prompt we used is given in the Appendix Table 14.
It is designed to ensure that the sequential order in which the responses are presented does not affect the judgement. The evaluation model is prompted to identify objects misaligned with the image context, identifying any discrepancies in count, position, and colors of objects in the images. The results presented in Table 7 show that SPIN improves Accuracy up to 7%, while maintaining Detailedness close to baseline. We use the same values for Layers, , and reported for CHAIR for all captioning tasks.
4.6 General Performance Assessment
MME Evaluation: In Table 5, we present SPIN results on MME full set Fu et al. (2023) that assesses the perceptual and cognitive abilities of VLMs across a total of 14 subtasks, including tasks such as OCR, visual knowledge, attribute relationships, and object recognition. SPIN outperforms the baseline on both perception and cognition tasks for all three models, with a small degradation in cognition for Qwen2.5-VL. The improvements in the first four subtasks relating to existence, count, position, and colour demonstrates the model’s ability to reduce object-level and attribute-level hallucinations.
MMMU Evaluation: In Table 6, we present MMMU Yue et al. (2024) (a standard multimodal understanding and reasoning benchmark) scores for SPIN to assess its impact on the general capability of LVLMs. SPIN attains the same performance as baseline for LLaVA-1.5 (7B) and Qwen2.5-VL, while attaining a small improvement on LLaVA-1.5 (13B) and LLaVA-Next. This demonstrates that head suppression does not hinder, but enhances the model’s core capabilities while providing substantial hallucination reduction. Notably, we use the same parameters for head suppression (Layers, , and ) as reported for CHAIR evaluation for both MME and MMMU.
LLaVA-1.5 (7B)
Layers
CS
CI
F1
132
0.05
17.4
5.8
74.6
132
0.10
22.8
7.4
74.1
132
0.15
20.8
6.4
73.8
116
0.05
18.2
6.2
73.7
120
0.05
18.4
6.3
74.1
124
0.05
19.2
6.5
73.8
432
0.05
42.0
11.4
78.2
832
0.05
42.4
11.1
78.8
LLaVA-1.5 (13B)
Layers
CS
CI
F1
132
0.05
45.8
12.0
78.2
132
0.10
29.6
8.0
78.9
132
0.15
41.6
10.3
80.2
132
0.20
36.0
10.0
79.1
116
0.10
30.6
8.3
79.6
120
0.10
29.2
7.9
79.1
124
0.10
28.0
9.2
79.3
Qwen-VL
Layers
CS
CI
F1
132
0.05
44.4
11.7
77.6
132
0.10
39.8
10.9
77.1
132
0.15
41.6
11.7
77.5
132
0.20
39.6
11.2
76.9
132
0.25
36.4
9.7
77.3
132
0.30
26.0
9.1
75.6
120
0.30
27.6
7.9
76.1
124
0.30
28.4
8.3
76.2
832
0.30
32.4
9.4
76.3
Table 8: Ablation on the ratio of suppressed heads and layer selection for three different models. The best is identified in the top part of the tables (marked with blue). The layer configuration for selected value of is presented in the bottom part (best layer configuration is marked in yellow).
Figure 5: Ablation on the suppression factor () for four models.
@@ 4.7 Throughput Estimation
To assess the efficiency of different hallucination mitigation strategies, we compare their throughput in Figure 6. Throughput is computed by measuring the total computational latency
for generating tokens, and dividing the total number of generated tokens by this latency. SPIN achieves the highest throughput among SoTA approaches, while achieving throughput performance close to baseline.
Figure 6: Throughput comparison with existing methods given by the number of tokens generated per second.
@@ 4.8 Ablation Studies
We present the hyperparameter evaluations for SPIN in Table 8, where we first vary to obtain reduced hallucinations without a significant drop in F1. We obtain for LLaVA 7B, 0.10 for LLaVA 13B, and 0.30 for Qwen-VL. Next, we evaluate different layer configurations to find hallucination-prone layers
. While uniformly pruning across all layers works best for LLaVA 7B, initial layers are more hallucination-prone for LLaVA 13B and Qwen-VL. Finally, for selected and layers for head suppression, we plot , and F1 for different in Figure 5. F1 scores show comparable values for different for MiniGPT-4 and LLaVA 13B models. However, for LLaVA 7B and Qwen-VL, complete pruning leads to a drop in F1.
This happens because the pre-trained model weights are not adapted to this change. Therefore, we perform partial head suppression using the parameter , where a higher value of preserves the F1 score by reducing the suppression intensity.
@@ 5 Conclusions
We propose an inference-only solution for ease of integration and to accommodate scenarios with limited training resources. However, our approach could be enhanced by incorporating a trainable router for more adaptive head selection. Our method’s effectiveness is somewhat reduced when paired with sampling-based decoding strategies (e.g., nucleus sampling). We attribute this to the greater token selection variability inherent in these stochastic approaches (as further detailed in Section 4).
Our method requires access to the model weights, which limits its applicability to open-source models. It cannot be applied to closed-source or API-based models, where direct access to weights is restricted.
@@ A.2 Ablation Studies
Head Selection Strategy:
We explore a variety of training-free head selection strategies, as summarized in Table 9. Query-based head selection proposed in Jin et al. (2024) is computed based on the L2 norm of the query tokens, whereas the key based selection is performed based on the L2 norm of the key tokens. In addition, we experiment with head selection guided by the overall attention (to both text and image tokens) as well as attention to the image tokens only. Among these strategies, selecting heads based solely on image attention proves to be the most effective for hallucination reduction, yielding the best CHAIR scores. Here, we use , completely pruning attention heads across all layers.
Selection
Supp.
CHAIRS
CHAIRI
F1
Strategy
Heads
Query
0.10
40.4
11.6
77.4
Query
0.25
34.6
10.4
75.8
Key
0.25
44.6
12.4
77.1
Attention
0.15
20.4
8.0
73.4
Attention
0.05
17.6
6.1
74.5
Image Attention
0.05
17.4
5.8
74.6
Table 9: Comparison of different training-free head selection strategies in SPIN on LLaVA-1.5 (7B) using CHAIR evaluation. Selecting heads based on image attention yields the most effective reduction in hallucination.
@@ A.3 Head Mask Visualization
The head masks for POPE and CHAIR using 95% active heads
are visualized in Figure 7. The pruned heads are more consistent for POPE, likely due to the fact that “Yes” or “No” answers are generated as the first token. In contrast, CHAIR generates longer sequences, causing tokens at each position to likely prune different heads.
Figure 7: Heatmap for head masks averaged over all test samples for CHAIR (left) and POPE random split (right) for 95% active heads. White and dark blue indicate the always pruned and always active heads, respectively.
Model
Mode
Layers
Accuracy
F1
LLaVA-1.5
Baseline
85.69
84.28
SPIN
116
0.15
0.0
87.84
87.13
SPIN
116
0.15
0.001
87.79
87.07
SPIN
132
0.20
0.08
88.19
87.42
LLaVA-1.5
Baseline
85.12
84.48
SPIN
132
0.35
0.0
88.47
88.00
SPIN
124
0.35
0.0
88.40
88.01
SPIN
120
0.35
0.001
88.50
88.17
MiniGPT-4
Baseline
78.08
76.20
SPIN
132
0.05
0.0
79.43
78.78
SPIN
124
0.05
0.08
79.37
78.61
SPIN
124
0.05
0.001
79.14
78.46
Shikra
Baseline
78.27
78.89
SPIN
116
0.20
0.0
79.20
79.40
SPIN
116
0.20
0.01
79.17
79.37
SPIN
116
0.20
0.05
79.19
79.45
Table 10: Multi-turn POPE evaluation under greedy decoding. SPIN consistently reduces hallucinations across all models on multi-turn vision-language tasks.